Whats the distance between the points (-5, -10) and (2, 4).

Answer by Boreal(15235)

distance is sqrt( distance between x points^2+distance between y points^2)
distance between -5 and 2 is 7 (don't worry about sign since it will be squared and become positive regardless)
7^2=49
distance between -10 and 4 is 14 and 14^2=196
The distance is sqrt (196+49)=sqrt (245)
This is sqrt(49)*sqrt(5)=7 sqrt(5) units.